# Domain Generalization in LiDAR Semantic Segmentation Leveraged by Density Discriminative Feature Embedding

### v1

@@ Abstract
While significant progress has been achieved in LiDAR-based perception, domain generalization continues to present challenges, often resulting in reduced performance when encountering unfamiliar datasets due to domain discrepancies. One of the primary hurdles stems from the variability of LiDAR sensors, leading to inconsistencies in point cloud density distribution. Such inconsistencies can undermine the effectiveness of perception models. We address this challenge by introducing a new approach that acknowledges a fundamental characteristic of LiDAR: the variation in point density due to the distance from the LiDAR to the scene, and the number of beams relative to the field of view. Understanding this, we view each LiDAR’s point cloud at various distances as having distinct density distributions, which can be consistent across different LiDAR models. With this insight, we propose the Density Discriminative Feature Embedding (DDFE) module, crafted to specifically extract features related to density while ensuring domain invariance across different LiDAR sensors. In addition, we introduce a straightforward but effective density augmentation technique, designed to broaden the density spectrum and enhance the capabilities of the DDFE. The proposed DDFE stands out as a versatile and lightweight domain generalization module. It can be seamlessly integrated into various 3D backbone networks, consistently outperforming existing state-of-the-art domain generalization approaches. We commit to releasing the source code publicly to foster community collaboration and advancement.
Index Terms:
Autonomous vehicles, 3D semantic segmentation, Domain Generalization, 3D LiDAR.I Introduction
@@
Light Detection and Ranging (LiDAR) provides detailed 3D data, making it indispensable for environmental perception in autonomous vehicles. Among the various LiDAR-based perception tasks, semantic segmentation plays a crucial role in understanding the driving scene by classifying each point into multiple classes. While LiDAR-based semantic segmentation [1, 2, 3] have been widely studied, their impressive performance is often constrained to scenarios where source and target datasets align perfectly. However, mismatches between these datasets can lead to significant performance declines. These are mainly attributed to two primary factors: variations in the environment — illustrated by datasets Waymo in the USA [4], KITTI in the Germany [5], and nuScenes in the Singapore [6]) — and sensor-induced discrepancies in the number of beams and Field of View (FoV). Numerous unsupervised domain adaptation studies [7, 8, 9] offer solutions but demand separate fine-tuning before deployment. In contrast, models trained with robust domain generalization methods hold greater promise for real-time autonomous driving systems. However, the crucial field remains under-researched and its potential is yet to be fully realized. Existing study [10] has pinpointed the point cloud density distribution as a prime performance hindrance. While some solutions like point cloud sampling and completion-based methods attempt to remedy this, they fall short in achieving the desired performance [11] and require sequentially labeled data and knowledge of ego-motion [12].
@@
In this paper, we introduce a new perspective in domain generalization, focusing on the challenges posed by density variations from different LiDAR sensors. Previous research [10, 13, 14] primarily concentrated on global density differences, typically considering datasets from 64-channel LiDARs like Waymo [4] denser than those from 32-channel LiDARs like nuScenes [6]. However, as shown in Fig. 1, density comparisons can be more nuanced. For example, a vehicle at 9 meters in Waymo appears denser than one at the same distance in nuScenes, but at different distances, such as a vehicle at 35 meters in Waymo versus 9 meters in nuScenes, the latter may appear denser. This inconsistency arises from an inherent property of LiDAR sensors where density varies with distance. Recognizing this, we view each LiDAR’s point cloud at various distances as having distinct density distributions, which can be consistent across different LiDAR models. This implies that despite the resolution differences in LiDAR sensors, overlapping density domains can exist within these multiple domains. For example, vehicles at certain distances in Waymo and nuScenes could have similar densities. This understanding of overlapping density domains provides vital insights for domain generalization.
@@
Building on these observations, we present a domain alignment method expertly, tailored to both discern and enhance shared density distributions. We hypothesize that even if the density distributions of source and unseen data differ, there may still be overlapping density regions between them. At the core of our strategy is the density-aware embedding module that modulates features by utilizing attention metrics, guided by our beam density estimation. The beam density estimation module is designed to encode the density inherent to each 3D voxel, facilitating a detailed differentiation of densities specific to certain regions. Furthermore, we introduce a density soft clipping technique. This method constrains the density spectrum, ensuring it does not encompass density distributions from unseen domains that are absent in source datasets. To supplement this, we use density augmentation to widen the density spectrum of the source data, thereby enhancing its domain generalization capabilities. Extensive experiments validate the superiority of our method over the conventional domain adaptation and generalization methods across multiple 3D backbone networks.
@@
We introduce a new perspective for domain generalization to overcome density variations caused by different LiDAR sensors, utilizing the characteristic of LiDAR where density varies with distance.
We propose a density discriminative feature embedding module, DDFE, designed to identify shared density distribution and subsequently amplify relevant features.
@@
Extensive experiments demonstrate that our method outperforms state-of-the-art Domain Generalization (DG), as well as Domain Adaptation (DA) approaches.
II Related Work
II-A LiDAR-based Semantic Segmentation
@@
LiDAR point clouds pose unique challenges due to their irregular, unordered, and unstructured nature, making standard image-based convolution operations inapplicable. Research on representing point cloud data can be broadly categorized into Projection-based, point-based, and voxel-based method. Projection-based methods transform a 3D point cloud into a 2D range view through either a spherical or bird’s-eye view projection. By doing so, they can utilize lightweight models like 2D convolution [15, 16, 17, 18] or transformer [3, 19] rather than 3D convolution. Nevertheless, such methods face the inherent limitation of 2D kernels not preserving the receptive field in the actual 3D space. Point-based methods [20, 21, 22] directly extract features from the point cloud. These methods utilize all the 3D spatial information without distortion. However, the trade-off is a substantial demand on memory and computational resources. In voxel-based approaches, point clouds are segmented into 3D grids, known as voxels, and features are extracted using 3D convolutional methods. However, the high computational and memory demands of dense 3D convolution have prompted recent research [23, 24, 1, 2] to adopt sparse 3D convolution techniques. Furthermore, studies [1, 2] have employed cylindrical partitioning techniques to effectively manage the challenges posed by uneven and imbalanced density distributions that occur at varying distances.
II-B LiDAR-based Domain Adaptation
@@
Differences in the LiDAR sensor or changes in the operating environment induce variations in the distribution of point cloud data. These variations can subsequently compromise the perception performance of autonomous driving systems. Constructing training datasets for every individual sensor and location to mitigate this degradation in performance is prohibitively expensive. In the quest for cost-effective performance maintenance, the spotlight has shifted towards unsupervised domain adaptation [25, 26, 7, 9, 27]. This strategy aims to leverage both the labeled source datasets and unlabeled target datasets to maintain the performance in the target domain. Yi et al. [11] leverage sequential LiDAR frame data to train a network that reconstructs dense frames, performing semantic segmentation in a canonical domain for adaptation. Saltori et al. [28] propose a source-free unsupervised domain adaptation (SF-UDA) strategy for 3D object detection, allowing model adaptation to target datasets without source dataset reference. Rochan et al. [29] propose a range view-based unsupervised domain adaptation method, aligning beam positions between training and target data, thereby ensuring spatial consistency in sensor data for model adaptation. Yuan et al. [8] present a multi-scale feature extraction technique to handle LiDAR density variations and propose an adversarial learning weight mechanism, capitalizing on consistency in class predictions across point and voxel representations. Shaban et al. [30] introduce LiDAR-UDA, combining point cloud subsampling to mimic various LiDAR types and refining pseudo labels through Cross-frame Ensembling and a Learned Aggregation Model (LAM). Although these domain adaptation methods demonstrate effectiveness, they share a common limitation: the need for individual fine-tuning with target data upon each domain shift.
II-C LiDAR-based Domain Generalization
@@
Domain Generalization aims to improve the performance in scenarios where the target domain data, unseen during training, is encountered. Research on domain generalization for LiDAR-based perception has recently witnessed a surge, as highlighted in a number of studies [31, 32, 10, 12, 33, 34, 35]. Kim et al. [10] emphasize generalizable representation learning. Their method leverages sparsity invariance feature consistency and bridges the semantic correlation consistency between the source data and the sparse domain generated by beam sampling. Ryu et al. [12] introduce a method to aggregate multiple frames to generate dense world models and augment data by sampling through randomized LiDAR configurations. However, this approach is limited in requiring knowledge of ego-motion and the sequentially labeled data. Li et al.[34] introduce BEV-DG to improve domain generalization in cross-modal semantic segmentation. They employ a bird’s-eye view for enhanced cross-modal learning and develop Density-maintained Vector Modeling (DVM) for efficient learning of domain-invariant features. Saltori et al.[35] explore domain generalization in LiDAR semantic segmentation, using both synthetic and real-world datasets. Their approach, LiDOG, utilizes semantic priors in a 2D bird’s-eye view to extract domain-agnostic features. Sanchez et al. [33] propose a label propagation approach incorporating multi-frame aggregation with known ego-motion. Nevertheless, this method is limited when the ego-motion is unknown and cannot run in real-time because it is based on KPConv [20], which is computationally expensive. Distinctively, our approach advances a single-frame domain generalization approach that considers the inherent characteristics of LiDAR. Consequently, our method offers the distinct advantage of bypassing the need for ego-motion and sequentially labeled data during the training and inference phases.
III Method
In this section, we introduce a domain generalization method for LiDAR semantic segmentation that leverages the unique characteristics of each LiDAR sensor, as shown in Tab. I and Fig. 2. First, we describe our point-voxel feature encoding network, adept at processing LiDAR point clouds by utilizing the capabilities of both point and voxel features. Subsequently, we detail our density embedding technique, designed to compute the expected density for each ray and modulate the encoded features to ensure generalization across different domains. Lastly, we introduce a simple yet effective data augmentation technique that extends the density range without the need for additional data, such as ego-motion or sequentially labeled data. The overall framework is shown in Fig. 3.
| Waymo | SemanticKITTI | nuScenes | ||||
| 64 | 64 | 32 | ||||
III-A Point-voxel feature encoding
@@
We introduce a technique to extract generalized representations across domains from point clouds. The distribution of intensity values varies depending on the LiDAR sensor. Leveraging intensity information can decrease the performance of perception models when transitioning between sensors. Thus, we omit the use of LiDAR intensity values to enhance domain generalization performance. Furthermore, each LiDAR sensor has a unique distribution of measurement errors, particularly within the intra-voxel region. To minimize variance induced by localized sensing noise, we incorporate a direct voxel-wise feature encoding into the PointNet-based feature extraction [36]. This direct encoding, which bypasses localized information within the voxels, effectively dismisses grid size variations, such as those of 20cm. Given a point cloud containing 3D points, we initially partition it into 3D voxels. For each voxel , it is defined as a set of points and is represented by , where indicates the number of points contained in each voxel. The coordinates of these points are subsequently transformed into the form of , where correspond to spherical coordinates. These transformed coordinates are then encoded into voxel-wise features through an MLP. We also encode point-wise features , which are essential to produce point-wise outputs. For each point in voxel centered at , we compute the offset as , with . Here, represents the number of points belonging to . These offsets are then processed by a point head to generate point-wise features.
III-B Beam density estimation module
@@
In this section, we detail the method to compute the density expectation for each ray emanating from a LiDAR sensor. Since all rays from a LiDAR originate from a singular source and are emitted at a fixed angle, the density associated with each beam can be deduced using a spherical projection informed by the beam configuration. The LiDAR sensor configuration, accessible from low-level sources such as an SDK, determines the set of horizontal and vertical inclinations and for each beam in this manner:
@@
where and denote the minimum and maximum LiDAR field of view, while and represent the number of horizontal and vertical beams, detailed in Tab. I.
To project LiDAR beam inclinations onto spherical projected image coordinates, we use the projection function defined as:
@@
where is the radial distance of a LiDAR point in the spherical coordinates. Taking into account that the density of the beam diminishes in proportion to the square of the distance due to its radial emission, we incorporate the inverse of in our computation. Moreover, if the LiDAR conducts uniform scans in the horizontal direction, the Gaussian filter of would yield a constant value across all outputs. In this case, we substitute with for computational efficiency.
III-C Density soft clipping
@@
When exposed to densities that deviate from those present in the source dataset, the model is susceptible to performance degradation in an unseen domain. To alleviate this issue, we introduce a density soft clipping method that confines the density spectrum using the function as follows:
@@
where is clipped density of point , and and are functions that extract and channel-wise percentile values from the density embedding of the training domain, respectively. Employing these percentile values allows us to discount the outlier density values. To calculate the percentile on the fly under memory constraints, we adopt the Reservoir Sampling technique with a sample size . A comprehensive description of the algorithm is provided in Alg. 1.
III-D Density-aware embedding module
@@
For each point , we derive a point-wise density embedding feature , using the beam density estimation module with density soft clipping. This feature serves as an input of both a point-wise attention function and a voxel-wise attention function , facilitating the generation of the domain-invariant density-discriminative features. In the case of point-wise attention, is dimensionally aligned with the point-wise feature to generate density discriminative feature as follows:
@@
where consists of two 1D convolution layers followed by a sigmoid function. For voxel-wise attention, we first compute the average of density embedding functions for all points within the voxel. Subsequently, the approach deployed for the point-wise feature is similarly applied, as follows:
@@
where is a clipped density of voxel and is a voxel-wise density discriminative feature and consists of two 1D convolution layers followed by a sigmoid function. An aggregation function for the voxel-wise feature is adjusted by voxel-wise attention and concatenated with the max pooled point-wise feature within the . The extracted features pass through one 1D convolutional layer and are then used as input to the 3D backbone.
| (a) Raw point cloud | (b) Augmented point cloud |
III-E Density Augmentation
Our density alignment method operates on the premise that various datasets include canonical density domains. However, when datasets differ significantly in their density ranges, this assumption may falter. To address this potential shortcoming, we introduce a data augmentation method designed to expand the density spectrum. Among current 3D point cloud augmentation techniques, several methods [37, 13, 10] aim to address density variations, either directly or indirectly. For instance, Mix3D [37] increases density by combining two point clouds, while beam drop augmentation [13, 10] decreases it by skipping certain LiDAR channels. However, these techniques offer limited density variation ranges. Specifically, Mix3D, which merges two point clouds based on centroid alignment without additional translation, typically results in densities of only one or two times the original level, falling short in achieving a broader spectrum of beam-density levels.
To overcome these limitations, we propose Enhanced-Mix3D, an improved approach described in Alg. 2. Our method merges points and densities from two randomly selected point clouds with additional random translation. This Enhanced-Mix3D not only introduces a more continuous range of density variations but also ensures higher entropy and increased data diversity. Moreover, it accounts for the ego vehicle’s trajectory in determining new overlapping centers, as depicted in Fig. 4, leading to a more varied and flexible augmentation strategy. The process of density aggregation is followed by density clipping, similar to the process used for a single sweep. Beyond the aforementioned modification, we also enhance the density spectrum using a beam drop augmentation technique. Inspired by the works of [13, 10], we sample the vertical beams of the LiDAR at intervals of 2, effectively reducing the point cloud density by half. During training, both the Enhanced-Mix3D and beam drop augmentations are set with a probability of 0.5.
IV Experiments

### v2

# Rethinking LiDAR Domain Generalization: Single Source as Multiple Density Domains

@@ Abstract
In the realm of LiDAR-based perception, significant strides have been made, yet domain generalization remains a substantial challenge. The performance often deteriorates when models are applied to unfamiliar datasets with different LiDAR sensors or deployed in new environments, primarily due to variations in point cloud density distributions. To tackle this challenge, we propose a Density Discriminative Feature Embedding (DDFE) module, capitalizing on the observation that a single source LiDAR point cloud encompasses a spectrum of densities. The DDFE module is meticulously designed to extract density-specific features within a single source domain, facilitating the recognition of objects sharing similar density characteristics across different LiDAR sensors. In addition, we introduce a simple yet effective density augmentation technique aimed at expanding the spectrum of density in source data, thereby enhancing the capabilities of the DDFE. Our DDFE stands out as a versatile and lightweight domain generalization module. It can be seamlessly integrated into various 3D backbone networks, where it has demonstrated superior performance over current state-of-the-art domain generalization methods. Code is available at https://github.com/dgist-cvlab/MultiDensityDG.
Keywords:
LiDAR Semantic Segmentation Domain Generalization
1 Introduction
@@
Light Detection and Ranging (LiDAR) provides detailed 3D data, making it indispensable for environmental perception in autonomous vehicles. Among the various LiDAR-based perception tasks, semantic segmentation plays a crucial role in understanding the driving scene by classifying each point into multiple classes. While LiDAR-based semantic segmentation [48, 13, 1] have been widely studied, their impressive performance is often constrained to scenarios where source and target datasets align perfectly. However, mismatches between these datasets can lead to significant performance declines. These are mainly attributed to two primary factors: environmental variations as exemplified by the Waymo [35] in the USA, the SemanticKITTI [2] in Germany, and the nuScenes [4] in Singapore; and sensor-induced discrepancies, including differences in the number of beams and the field of view (FOV).
Numerous Unsupervised Domain Adaptation (UDA) studies [30, 45, 17] address these issues but require additional fine-tuning whenever the target domain changes. In contrast, models deployed with robust Domain Generalization (DG) techniques are increasingly sought after for their potential to enable real-time autonomous driving systems without needing constant fine-tuning. Nevertheless, this crucial field remains under-researched, and its potential has yet to be fully realized. Existing study [16] pinpoints the point cloud density distribution as a prime performance hindrance. While some solutions like point cloud sampling and completion-based methods attempt to remedy this, they fall short in achieving the desired performance [44] and require sequentially labeled data and knowledge of ego-motion [29].
@@
In this paper, we introduce a novel perspective in domain generalization by addressing the challenges posed by density variations across different LiDAR sensors. Previous studies [16, 38, 14] have primarily focused on global density differences, typically considering datasets from 64-channel LiDARs like Waymo [35] denser than those from 32-channel LiDARs like nuScenes [4]. However, these approaches oversimplify the inherent complexity of LiDAR data, which exhibits a wide range of density spectra over distance, and interpret them as a single global density value. Contrary to the simplified view of global density comparisons, our research recognizes that LiDAR point clouds are composed of regions with varying densities, as shown in Fig. 1. These variations are influenced by the distance of objects from the LiDAR sensor, leading to a more complex and nuanced understanding of density within LiDAR data.
@@
To do so, we propose a Density Discriminative Feature Embedding (DDFE) module, designed to enhance the domain generalization capabilities of LiDAR-based segmentation networks by exploiting these density variations. The DDFE module incorporates a beam density estimation module that encodes the specific densities for each 3D voxel, enabling refined density discrimination across regions. The features are then modulated using attention mechanisms guided by beam density, further enhancing the model’s ability to generalize across domains with varying density characteristics. Furthermore, we introduce a density soft clipping technique. This method constrains the density spectrum, ensuring it does not encompass density distributions from unseen domains that are absent in source datasets. To supplement this, we use density augmentation to widen the density spectrum of the source data, thereby enhancing its domain generalization capabilities. Extensive experiments validate the superiority of our method over the conventional domain adaptation and generalization methods across multiple 3D backbone networks.
@@
We introduce a new perspective for domain generalization to overcome density variations caused by different LiDAR sensors, utilizing the diverse densities within point clouds observed in the source domain.
We propose a density discriminative feature embedding module designed to identify areas with similar density distribution and amplify relevant features.
@@
Extensive experiments demonstrate that our method outperforms state-of-the-art Domain Generalization and Domain Adaptation methods.
2 Related Work
2.1 LiDAR-based Semantic Segmentation
@@
LiDAR point clouds pose unique challenges due to their irregular, unordered, and unstructured nature. Consequently, approaches to represent point cloud data have been broadly classified into three categories: Projection-based, point-based, and voxel-based methods. Projection-based methods transform a 3D point cloud into a 2D representation through either a spherical or bird’s-eye view projection. By doing so, they can utilize lightweight models like 2D convolution [8, 47, 25, 5] or transformer [1, 6] rather than 3D convolution. Nevertheless, such methods face the inherent limitation of 2D kernels not preserving the 3D geometric information of the real world. Point-based methods [36, 15, 9] directly extract features from the point cloud. These methods utilize all the 3D spatial information without distortion. However, the trade-off is a substantial demand on memory and computational resources. In voxel-based approaches [11, 7, 48, 13], point clouds are quantized into 3D grids, known as voxels, and features are extracted using 3D convolutional methods. Alongside sparse convolution [12], voxel-based methods achieve high performance by maintaining an actual geometric receptive field while ensuring an efficient computational load.
2.2 LiDAR-based Domain Adaptation
@@
Variations in LiDAR sensor configurations, as shown in Tab. 1, or alterations in operating environments can lead to differences in the distribution of point cloud data, which, in turn, may adversely affect the perception capabilities of autonomous driving systems. To tackle this challenge without constructing new training datasets for every sensor type and location, research has increasingly focused on unsupervised domain adaptation techniques [39, 31, 43, 30, 45, 17, 26]. These methods strive to preserve performance in the target domain by utilizing labeled data from a source domain alongside unlabeled data from the target domain. C&L [44] leverages sequential data to train a voxel completion network and segments on the reconstructed canonical domain to overcome domain discrepancies. Rochan et al. [28] propose a range-view-based unsupervised domain adaptation method, aligning beam positions between training and target data. LiDAR-UDA [34] employs beam subsampling to mimic different LiDAR sensors, and utilizes cross-frame ensembling and a Learned Aggregation Model to acquire improved pseudo labels. Although these DA methods demonstrate effectiveness, they share a common limitation: the need for individual fine-tuning with target data upon each domain shift.
2.3 LiDAR-based Domain Generalization
@@
Domain Generalization [19, 41, 16, 29, 33, 21, 32] aims to improve the performance in scenarios where the target domain data, unseen during training, is encountered. DGLSS [16] leverages sparsity invariance feature consistency and bridges the semantic correlation consistency between the source data and the sparse domain generated by beam sampling. LiDomAug [29] aggregates multiple frames to generate dense world models and augment data by sampling through randomized LiDAR configurations with additional knowledge of ego-motion and the sequentially labeled data. BEV-DG [21] employs a bird’s-eye view for enhanced cross-modal learning and develops density-maintained vector modeling for efficient learning of domain-invariant features. LiDOG [32] utilizes semantic priors in a 2D bird’s-eye view to extract domain-agnostic features. Sanchez et al. [33] introduce a label propagation method that integrates multi-frame aggregation and ego-motion, although it demands significant computational resources due to the adaptation of KPConv [36]. Distinctively, our approach advances a single-frame domain generalization approach that considers the inherent characteristics of LiDAR. Consequently, our method offers the distinct advantage of bypassing the need for ego-motion and sequentially labeled data during the training and inference phases.
3 Method
In this section, we detail our domain generalization method for LiDAR semantic segmentation, which capitalizes on density variations observable within a single source LiDAR. The proposed Density Discriminative Feature Embedding (DDFE) module is primarily composed of four components: Point-voxel feature encoding as detailed in Sec. 3.1, Beam density estimation module in Sec. 3.2, Density soft clipping in Sec. 3.3, and Density-aware embedding module in Sec. 3.4. Additionally, Sec. 3.5 outlines our density augmentation technique. The overall framework of the proposed DDFE module is shown in Fig. 2.
3.1 Point-voxel feature encoding
@@
We introduce a technique to extract generalized representations across domains from point clouds. The intensity distribution in point clouds is influenced by the specific LiDAR sensor used, which can adversely affect the performance of perception models during sensor transitions due to the variance in intensity information. Thus, we omit the use of LiDAR intensity values to enhance domain generalization performance. Furthermore, each LiDAR sensor has a unique distribution of measurement errors, particularly within the intra-voxel region. To minimize variance induced by localized sensing noise, we incorporate a direct voxel-wise feature encoding into the PointNet-based feature extraction [27]. This direct encoding, which bypasses localized information within the voxels, effectively dismisses grid size variations, such as those of 20cm. Given a point cloud containing 3D points, we initially partition it into 3D voxels. For each voxel , it is defined as a set of points and is represented by , where indicates the number of points contained in each voxel. The coordinates of these points are subsequently transformed into the form of , where correspond to spherical coordinates. These transformed coordinates are then encoded into voxel-wise features through an MLP. We also encode point-wise features , which are essential to produce point-wise outputs. For each point in voxel centered at , we compute the offset as , with . Here, represents the number of points belonging to . These offsets are then processed by a point head to generate point-wise features.
3.2 Beam density estimation module
@@
In this section, we detail the method to compute the density expectation for each ray emitted from a LiDAR sensor. Since all rays from a LiDAR originate from a singular source and are emitted at a fixed angle, the density associated with each beam can be deduced using a spherical projection informed by the beam configuration. The LiDAR sensor configuration, accessible from low-level sources such as an SDK, determines the set of horizontal and vertical inclinations and for each beam in this manner:
@@
where and denote the minimum and maximum LiDAR field of view, while and represent the number of horizontal and vertical beams, detailed in Tab. 1. To project LiDAR beam inclinations onto spherical projected image coordinates, we use the projection function defined as:
@@
where is the radial distance of a LiDAR point in the spherical coordinates, and denotes the concatenation operation. Taking into account that the density of the beam diminishes in proportion to the square of the distance due to its radial emission, we incorporate the inverse of in our computation.
3.3 Density soft clipping
@@
When exposed to densities that deviate from those in the source dataset, the model is susceptible to performance degradation in an unseen domain. To alleviate this issue, we introduce a density soft clipping method that confines the density spectrum using the function as follows:
@@
where is clipped density of point , and and are functions that extract and channel-wise percentile values from the density embedding of the training domain, respectively. Employing these percentile values allows us to discount the outlier density values. To calculate the percentile on the fly under memory constraints, we adopt the Reservoir Sampling [37] technique with a sample size . Please refer to the supplementary material for a more detailed algorithm description.
3.4 Density-aware embedding module
@@
For each point , a point-wise density embedding feature is obtained via the beam density estimation module incorporating density soft clipping. This feature serves as an input of both a point-wise attention function and a voxel-wise attention function . These mechanisms are pivotal for crafting domain-invariant density-discriminative features. In the point-wise attention framework, is synchronized dimensionally with the point-wise feature , facilitating the creation of a density discriminative feature as follows:
@@
where consists of two 1D convolution layers followed by a sigmoid function. For voxel-wise attention, it begins by averaging the density embedding features for all points encapsulated within a voxel. This aggregated measure then undergoes a similar treatment as follows:
@@
where is a clipped density of voxel and is a voxel-wise density discriminative feature and consists of two 1D convolution layers followed by a sigmoid function. An aggregation function for the voxel-wise feature is adjusted by voxel-wise attention and concatenated with the max pooled point-wise feature within the . Finally, the extracted features are passed through a single 1D convolutional layer, resulting in 32-channel features. These density-aware features are subsequently fed into a 3D backbone for LiDAR semantic segmentation.
3.5 Density Augmentation
The DDFE module aims to align densities across varied sensor domains by considering each dataset as a collection of multiple density domains. This method, while effective in many scenarios, may face challenges when there’s little to no overlap in the density ranges between the source and unseen datasets. To address this potential issue, we introduce a density augmentation method aimed at broadening the density spectrum covered by the training data. Several existing 3D point cloud augmentation strategies, such as [22, 38, 16] outlined in recent studies, attempt to address density variations either directly or by implication. We adapt the Mix3D [22] for our purposes, which we refer to as enhanced-Mix3D, incorporating random translations along the direction of ego-vehicle movement, along with additional rotational transformations to simulate a wider range of density variations. We also employ beam sampling technique [38, 16], selectively eliminating specific LiDAR beams, to amplify the density in the lower direction. During the training phase, we apply both enhanced-Mix3D and beam sampling augmentations with a set probability of 0.5, aiming to effectively simulate diverse density conditions. Please refer to the supplementary material for more detailed implementation.
4 Experiments